OpenStudy (anonymous):

Please help me solve this and check for extraneous solutions. 10 over x+4=15 over 4x+4

OpenStudy (anonymous):

@oaktree

OpenStudy (oaktree):

Right here.

OpenStudy (oaktree):

Is it this:\[\frac{ 10 }{ x+4 } = \frac{ 15 }{ 4x+4 }\]

OpenStudy (anonymous):

Yay, yes

OpenStudy (anonymous):

@oaktree ?

OpenStudy (anonymous):

"cross multiply" and get \[\frac{ 10 }{ x+4 } = \frac{ 15 }{ 4x+4 }\iff 10(4x+4)=15(x+4)\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Okay 40x+40=15x+60

OpenStudy (anonymous):

yeah three more steps

OpenStudy (anonymous):

Yes I believe so. Thank you but what would I do with a problem like....x over x-2 plus 1over x-4=2 over x squared-6x+8

OpenStudy (anonymous):

It has 3 instead of two @satellite73

OpenStudy (anonymous):

i would try to write it so i can understand it is it \[\frac{x}{x-2}+\frac{1}{x-4}=\frac{2}{x^2-6x+8}\]

OpenStudy (anonymous):

\[\frac{x}{x-2}+\frac{1}{x-4}=\frac{3}{x^2-6x+8}\]

OpenStudy (anonymous):

that right?

OpenStudy (anonymous):

It is the first one

OpenStudy (anonymous):

add on the left and get \[\frac{x}{x-2}+\frac{1}{x-4}=\frac{2}{x^2-6x+8}\] \[\frac{x(x-4)+(x-2)}{(x-2)(x-4)}=\frac{2}{(x-2)(x-4)}\]

OpenStudy (anonymous):

cooked up so that the denominator on the left is the same as the denominator on the right after you factor

OpenStudy (anonymous):

since the denominators are the same, you can equate \[x(x-4)+x-2=2\]and solve the quadratic

OpenStudy (anonymous):

ooooh you have to be real real careful here

OpenStudy (anonymous):

when you solve the quadratic, one solution will be \(x=4\) but that is not a solution to the original equation, because the original expressions are undefined if \(x=4\) so omit that solution

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

(Omit)

OpenStudy (anonymous):

what does what mean?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh, it means don't put that as a solution, because it is not one

OpenStudy (anonymous):

The answer that I got is (x+1)(x+2)

OpenStudy (anonymous):

Oh, thankyou

OpenStudy (anonymous):

Is that wrong?

OpenStudy (anonymous):

\[x(x-4)+x-2=2\] \[x^2-4x+x-2=2\] \[x^2-3x-4=0\] \[(x-4)(x+1)=0\]

OpenStudy (anonymous):

solutions to this are \(x=4\) or \(x=-1\) but as i said, \(4\) is not a solution to the original equation as it would make the denominator zero

OpenStudy (anonymous):

So then would I put both down answers?

OpenStudy (anonymous):

no you put down \(x=-1\) only

OpenStudy (anonymous):

Can you help me with one more that is different from that one?

OpenStudy (anonymous):

@satellite73